OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 1. 11 2. 19 3. 16 4. 8

OpenStudy (anonymous):

1

OpenStudy (lalaly):

g(4)=(4)^2+3=16+3=19 so the answer is 2